A nuclear fission reaction is defined to be a process in which a large nucleus divides into two

roughly equal parts.

(i) What is the potential at a distance of 4.50×10–14 m from one of the parts that has 53 protons in it?

(ii) What is the potential energy in eV of a similarly charged part at this distance?


1
Expert's answer
2021-01-22T11:27:12-0500

The electric potential due to protons at distance r is

"V=\frac{kQ}{r}"

"=\frac{(9\times 10^9 N.m^2/C^2)(53\times1.60\times10^{-19}C)}{4.50×10^{–14}}"

"=1.696 \times 10^6 V"

The electrostatic potential energy is

"U=\frac{kQ^2}{r}"

"=\frac{(9\times 10^9 N.m^2/C^2)(53\times1.60\times10^{-19}C)^2}{4.50×10^{–14}m}"

"=1.438 \times 10^{-11} J"